A toy manufacturer needs a piece of plastic in the shape of a right triangle with the longer leg 22 cm more than the shorter leg and the hypotenuse 44 cm more than the shorter leg. How long should the sides of the triangle​ be?
Answer by addingup(3677)

Let x be the short leg:
x^2+(x+22)^2 = (x+44)^2
2x^2+44x+484 = x^2+88 x+1936
Subtract x^2+88 x+1936 from both sides:
x^2-44x-1452 = 0
(x-66)(x+22) = 0
x-66 = 0 or x+22 = 0
x = 66 or x = -22
x = 66 or x = -22
We are not looking for a negative number, so let's try 66:
:
Short leg: 66
Long leg: 66+22 = 88
Hypotenuse: 66+44 = 110
Check:
a^2+b^2 = c^2
66^2+88^2 = 110^2
4356+7744 = 12100 Correct